find the measure of the reference angle for 400 degrees
Answer by drj(1380)

Find the measure of the reference angle for 400 degrees

The angle should be less than 360 degrees since a full circle is 360 degrees.
So 400-360=60 degrees.

Another example let's say we have an angle of 145 degrees, then its reference angle is 180-145=45 degrees. So reference angles need to be less than 90 degrees.

Another example let's say we have an angle of 215 degrees. Then its reference angle is 215-180=45 degrees.

I hope the above steps were helpful.

For FREE Step-By-Step videos in Introduction to Algebra, please visit http://www.FreedomUniversity.TV/courses/IntroAlgebra and for Trigonometry visit http://www.FreedomUniversity.TV/courses/Trigonometry.

And good luck in your studies!

Respectfully,
Dr J
http://www.FreedomUniversity.TV